# INDIC QA BENCHMARK: A Multilingual Benchmark to Evaluate Question Answering capability of LLMs for Indic Languages

### v1

@@ 3.2 Data Curation Methodology
Ext Gen Ext Gen Ext Gen Ext Gen
As 13.86 5.04 21.70 4.92 19.23 3.85 1.14 0.02
Bn 17.84 7.88 21.15 4.93 19.14 4.22 1.49 0.03
Gu 13.27 7.15 24.90 6.36 20.27 5.35 4.07 0.05
Hi 21.69 7.65 34.37 5.83 41.53 8.08 11.83 0.004
Kn 15.63 7.40 24.47 5.63 20.39 4.52 0.31 0.00
Ml 19.22 8.36 25.20 5.20 22.80 5.83 0.68 0.002
Mr 15.12 6.69 23.96 5.43 36.12 7.56 2.07 0.00
Od 11.11 5.88 9.06 2.58 11.23 5.31 0.26 0.00
Pa 15.60 6.87 28.43 6.28 21.96 5.92 1.31 0.00
Ta 19.96 9.07 22.45 5.30 19.74 4.99 0.70 0.00
Te 16.07 7.89 23.93 7.32 13.59 3.17 0.53 0.00
Average 16.31 7.26 23.60 5.43 22.36 5.35 2.22 0.0096
@@
Table 2: Performance of various Large Language Models on Zero-Shot INDIC QA BENCHMARK. We report the average numbers across span-extraction datasets and abstractive question-answering datasets.
@@
Table 3: Indic-RAG Dataset Statistics. Indic-RAG benchmark is a compilation of existing datasets, English datasets translated to Indian languages and Synthetic dataset generated using Gemini
Languages Bloom Gemma Llama-3 Openhathi
1 shot 3 shot 1 shot 3 shot 1 shot 3 shot 1 shot 3 shot
As 21.22 22.44 37.91 40.43 28.61 29.42 3.15 3.14
Bn 19.82 26.90 42.64 37.85 32.71 24.95 6.86 6.56
Gu 20.17 22.16 44.26 47.40 26.76 22.46 3.65 4.61
Hi 34.08 36.96 54.95 56.95 54.79 57.11 10.27 22.1
Kn 19.99 19.93 37.74 35.80 22.76 15.60 0.33 2.26
Ml 25.82 23.84 42.08 31.10 28.94 22.60 6.59 5.50
Mr 26.35 25.99 47.08 37.39 47.17 40.7 5.82 8.63
Od 15.72 15.68 16.66 22.05 13.37 5.39 0.37 3.32
Pa 23.18 23.98 42.18 47.17 31.03 30.62 2.00 3.52
Ta 26.80 26.97 43.27 43.90 28.66 27.06 4.51 4.94
Te 20.83 25.22 39.90 43.62 25.37 20.86 3.09 2.76
@@
Table 4: Performance of various Large Language Models on few-Shot INDIC QA BENCHMARK. We report the average numbers across span-extraction datasets and question-answering datasets.
4 Experiment Setup
We conducted a series of experiments to evaluate existing LLM’s performance, utilizing the NVIDIA RTX A100 both 40Gb and 80Gb variants for our computational needs. Our computational needs signify GPUs both for Translation and evaluation over the models. For inference or evaluation we utilized VLLM Kwon et al. (2023) which is an open-source library that supports LLM inference efficiently.
We evaluate the following LLMs on our benchmark: Open Hathi and its Instruction Finetuned variant (IFV) known as Airavata Gala et al. (2024), Bloom Le Scao et al. (2022) and its IFV named Bloomz, GemmaTeam et al. (2024), and its instruction fine-tuned variant Gemma-IT. Open Hathi [^3] https://www.sarvam.ai/blog/announcing-openhathi-series (7B parameter model), which was created through continual pre-training on the LLaMA-2 model Touvron et al. (2023). Airavata Gala et al. (2024) (7B parameter model) is an instruction fine-tuned version of OpenHathi. Both OpenHathi and Airavata are specifically trained for Hindi.
Gemma and Gemma-IT (7B parameter models) [^4] https://ai.google.dev/gemma/docs, which were released by Google. Though these models are not specifically trained for Indian languages, they exhibit multilingual capability. Aya-8B Aryabumi et al. (2024) is another instruction-tuned model designed explicitly for multilingual languages. LLaMA 3 and LLaMA 3 Instruct models  [^5] https://ai.meta.com/blog/meta-llama-3/(8B parameter model), part of the LLaMA family. Llama-3 has seen data from around 30 languages excluding English. This diverse set of models allows for a comprehensive evaluation of the strength and performance of our benchmarks across different architectures and training methods.
4.1 Evaluation Metrics
We borrow the key metrics of both Extractive and generative Question Answering as mentioned below:
1. F1 (macro-averaged) score The F1 score is a metric that represents the harmonic mean of precision and recall. The F1 score calculates the average similarity between predicted and actual answers by comparing the sets of words or tokens in the predicted and ground truth sentences.
2. Exact Match The Exact Match metric computes the percentage of instances that exactly match the ground truth. This metric is more stringent but evaluates the thoroughness of the model, as it results in either a true or false outcome for individual instances.
3. ROUGE(L) A Recall-Oriented Understudy for Gisting Evaluation metric is mostly used for evaluating summarization, we have used it to evaluate generative QA tasks.
5 Results and Analysis
On the other hand, OpenHathi was specifically trained on the Hindi language and so is its instruction finetuning variant Airavata. As a result, the performance of OpenHathi is significantly poor in all languages. Airavata benefits from further instruction fine-tuning on Hindi data and improves over OpenHathi for Hindi language but suffers poorly for other Indian languages. Bloomz produces the highest jump compared to Bloom and we hypothesize this is because a good portion of evaluation benchmark coming from generic-domain such wikipedia data has been seen by Bloomz during its training and instruction finetuning, making it a good choice for applications which aims to use common world knowledge.
Extractive vs Abstractive tasks: While it is clear that instruction finetuning helps more in abstractive QA tasks, both Table 2 and Table 3 shows positive correlation between the scores for extractive task and abstractive task across languages i.e. whenever numbers have improved for extractive QA tasks, it also improved for abstractive QA too. This is almost true for all the base and instruct variants of the models except Gemma, where Gemma instruct improves abstractive QA score but detoriates in extractive QA task. Careful analysis shows, generative task metrics change moderately between base models and their instruction finetuned variants. This is expected because generative metrics such as rogue(l) are more heuristic driven in nature and designed to ignore small variations, natural in non-deterministic text generation, unlike extractive metrics such as EM, F1. Thus generative metrics deviate in smaller scale than extractive metrics. However, the positive correlation between both the task metrics across models clearly establishes that the factors affecting the overall performance of the models shows similar signs for both extractive and generative tasks and hence improving one will likely improve the other as well.
On how to choose a model: Going by the results so far, one would pick BloomZ if the application needs only common world knowledge and needs a model which does well OOB. If there is use-case for which we have adequate Indic language finetuning data , it might be good to build over the world knowledge acquired by Gemma and Llama-3 and do instruction finetuning on Indic languages to make it better suitable for abstractive QA task. If we are very specific about a certain niche domain in only Hindi language, where common world knowledge is not a pre-requisite, Airavata can be a good candidate given its focus on Hindi based training and improvements in both extractive and abstractive task with instruction finetuning.
6 Conclusion
In this paper, we release a benchmark for evaluating the grounded Question-Answering capabilities of existing Large Language Models (LLM). The benchmark comprises datasets testing both extractive and abstractive answering abilities. Our observations reveal that question-answering capability in LLMs can be enhanced through instruction-tuning of these models with target language data. We hope that this benchmark will aid researchers in improving LLMs’ grounded question-answering abilities.
7 Limitations
While our research aims to provide a challenging and comprehensive benchmark for evaluating the LLMs on Hindi QA task, there might be some limitations. 1. Availability of high quality datasets for Hindi is limited. Despite our best efforts to curate the benchmark from various sources, it might still have an inherent bias introduced during the data collection/translation process. 2. Although we conducted quality checks, there might be subjective interpretability issues with the translated datasets. 3. While we attempted to diversify across various domains, the benchmark may not depict the true performance in a completely unseen domain.
Appendix A Appendix
A.1 Miscellaneous
@@
Figure 8: Base Model(Bloom) prediction on one of the instance of Generative QA dataset
Figure 9: Instruction tunned Model(Bloomz) prediction on one of the instance of Generative QA dataset.

### v2

@@ 3 Experiments
1 shot 3 shot 1 shot 3 shot 1 shot 3 shot 1 shot 3 shot
As 21.22 22.44 37.91 40.43 28.61 29.42 3.15 3.14
Bn 19.82 26.90 42.64 37.85 32.71 24.95 6.86 6.56
Gu 20.17 22.16 44.26 47.40 26.76 22.46 3.65 4.61
Hi 34.08 36.96 54.95 56.95 54.79 57.11 10.27 22.1
Kn 19.99 19.93 37.74 35.80 22.76 15.60 0.33 2.26
Ml 25.82 23.84 42.08 31.10 28.94 22.60 6.59 5.50
Mr 26.35 25.99 47.08 37.39 47.17 40.7 5.82 8.63
Od 15.72 15.68 16.66 22.05 13.37 5.39 0.37 3.32
Pa 23.18 23.98 42.18 47.17 31.03 30.62 2.00 3.52
Ta 26.80 26.97 43.27 43.90 28.66 27.06 4.51 4.94
Te 20.83 25.22 39.90 43.62 25.37 20.86 3.09 2.76
@@
Table 4: Performance of various Large Language Models on few-Shot Extractive INDIC QA BENCHMARK. We report the average (F1-score) across span-extraction datasets and question-answering datasets.
3.1 Evaluation Metrics
We chose widely used QA evaluation metrics for evaluating both extractive and abstractive Question Answering datasets:
1. F1 (macro-averaged) score: This score represents the harmonic mean of precision and recall, calculating the average similarity between predicted and actual answers by comparing the sets of words or tokens in the predicted and ground truth sentences.
2. ROUGE(L): A Recall-Oriented Understudy to measure how much of the words from the gold response are present in the generated response. This metric is commonly used for generative tasks such as summarization, and we have used it to evaluate abstractive QA tasks.
3.2 Translate-Test Inferences
In addition to direct inference, we conducted experiments on translate-test 4 inferences involving the following steps:
1. Use IndicTrans2 to translate the source language input (context and question) to English.
2. Prompt the LLM with the translated input and get the response.
3. Back-translate the generated LLM’s English response to the source language.
This method ensures the evaluation of the LLMs’ performance across languages by utilizing translation systems to bridge the gap between languages during inference.
4 Result and analysis
On the other hand, OpenHathi was specifically trained on the Hindi language and so is its instruction finetuning variant Airavata. As a result, the performance of OpenHathi is significantly poor in all languages. Airavata benefits from further instruction finetuning on Hindi data and improves over OpenHathi for Hindi language but suffers poorly for other Indian languages. Bloomz produces the highest jump compared to Bloom and we hypothesize this is because a good portion of evaluation benchmark coming from generic-domain such as Wikipedia data has been seen by Bloomz during its training and instruction finetuning, making it a good choice for applications which aims to use common world knowledge.
Extractive vs Abstractive Tasks: While it is clear that instruction finetuning helps more in abstractive QA tasks, both Table 5 and Table 6 show a positive correlation between the scores for extractive tasks and abstractive tasks across languages. This is almost true for all the base and instruct variants of the models except Gemma, where Gemma instruct improves the abstractive QA score but deteriorates in the extractive QA task. Careful analysis shows that abstractive task metrics change moderately between base models and their instruction finetuned variants. This is expected because abstractive metrics such as Rouge-L are more heuristic-driven in nature and designed to ignore small variations, natural in non-deterministic text generation, unlike extractive metrics such as F1. Thus abstractive metrics deviate on a smaller scale than extractive metrics. However, the positive correlation between both task metrics across models clearly establishes that the factors affecting the overall performance of the models show similar signs for both extractive and abstractive tasks and hence improving one will likely improve the other as well.
On How to Choose a Model: Going by the results so far, one would pick BloomZ if the application needs only common world knowledge and needs a model which does well out-of-the-box. If there is a use-case for which we have adequate Indic language finetuning data, it might be good to build over the world knowledge acquired by Gemma and Llama-3 and do instruction finetuning on Indic languages to make it better suitable for abstractive QA tasks. If we are very specific about a certain niche domain in only the Hindi language, where common world knowledge is not a pre-requisite, Airavata can be a good candidate given its focus on Hindi-based training and improvements in both extractive and abstractive tasks with instruction finetuning. However, the Aya models are particularly well-suited if we are specifically seeking high-quality instruction-tuned models for Indian languages.
Translate Test is an Effective Alternative to Source Language: Translation-based approaches are often more effective than direct generation in source languages using multilingual models. For languages like Punjabi, Gujarati, and Oriya, the translation-based approach outperforms the multilingual model. This comparison divides languages into two categories: (1) those where multilingual models perform better and (2) those where translation-based approaches are superior.
Mid-resource languages like Hindi and Bengali benefit more from multilingual models, while low-resource languages like Oriya and Punjabi perform better with translation-based approaches. This is because multilingual models struggle with insufficient language-specific data, leading to poor performance. In contrast, translation-based approaches leverage high-resource languages for reasoning and generation, making it easier to learn the translation task from multilingual data.
5 Conclusion
In this paper, we present a benchmark for evaluating the grounded Question-Answering (QA) capabilities of Large Language Models (LLMs) on both extractive and abstractive tasks. Our findings reveal that instruction-tuning with target language data significantly enhances QA performance, while the Translate-Test technique yields better results for low-resource languages. In contrast, high-resource languages benefit more from source language inference due to larger training datasets.
Despite advances in multilingual training, LLMs struggle with transfer learning to low-resource languages. Integrating effective translation systems into the QA pipeline is crucial for improving performance in these contexts. By releasing this benchmark, we aim to promote further research into the QA capabilities of LLMs across various languages, particularly for those that are underrepresented.
6 Limitation
Our research aims to provide a challenging and comprehensive benchmark for evaluating LLMs on the Hindi QA task, but it does face several limitations.
(1) The availability of high-quality datasets for Hindi is limited. Despite our best efforts to curate the benchmark from various sources, there might still be an inherent bias introduced during the data collection and translation process. Additionally, although we conducted quality checks, there may be subjective interpretability issues with the translated datasets. (2) while we attempted to diversify across various domains, the benchmark may not depict the true performance in a completely unseen domain.
Despite the strong performance of the Translate Test technique, particularly for low-resource languages, it has several limitations. One significant drawback is the potential for cascading errors. Translation errors occurring early in the pipeline can propagate through subsequent stages, adversely affecting the final output. This issue is typical in models that rely on sequential processing, where initial inaccuracies can compound over time. Moreover, while the Translate Test approach currently shows promising results, the ideal solution is the development of robust multilingual models that can handle and mitigate cascading errors effectively. Such models should be capable of generalizing well across various languages without relying heavily on error-prone translation processes. Our analysis focused on base models rather than instruction-tuned models. Previous research has shown that multilingual instruction-tuned A well-optimized multilingual instruction-tuned model could potentially address these limitations.
However, the challenge remains in developing effective instruction-tuning data for low-resource languages, whether through translation or other methods. This underscores the need for continued research to enhance instruction-tuning strategies in multilingual settings.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. This work is supported by and is part of BharatGen (https://bharatgen.tech/), an Indian Government-funded initiative focused on developing multimodal large language models for Indian languages.
Appendix A Appendix
A.1 Synthetic Data Creation
As discussed in 2.1, in addition to the translation dataset from high-resource to low-resource languages, we also generated data using the Gemini model Team et al. (2023). We chose this model due to its strong multilingual performance and considerable parameter count. To start, we sampled paragraphs from various sources, including Wikipedia, storybooks, Indian news articles, and a selection of books, aiming for domain diversity while incorporating Indian cultural nuances. We developed our benchmark in Hindi and selected a subset of the generated data for human verification. The verification process focused on several parameters, such as the fluency of the generated questions and answers, and whether the questions were directly derived from the paragraphs. Given that large language models can hallucinate even with clear instructions annotators were asked to score each instance on a scale of 1 to 5, where 1 indicated irrelevance and 5 indicated exact relevance. We also evaluated whether the answers were generated from the text or relied on external knowledge.
Once verified, we translated these data instances into other Indian languages. The synthetic data primarily consisted of texts from Ramayana textbooks, recent Indian news articles, cultural articles from Wikipedia, and classic storybooks, Dharmpal books. Most of the data was extracted from books using Optical Character Recognition (OCR), followed by a thorough cleaning process to compile useful passages.
A.2 Analysis
Indic Languages are Medium or Low Resource: Before diving into the results, it’s important to consider the resource availability in Indic languages. This influences the best strategy for QA tasks in these languages. Figure 2 shows the approximate number of tokens available for each language from sources like Wikipedia, websites, and PDFs. Compared to English, which has trillions of tokens, Indic languages have far fewer resources. Languages like Hindi and Bengali are medium-resource, while most others are low-resource, with some like Oriya and Punjabi being very low-resource. Although the statistics in Figure 2 are estimates and change over time and across different models, classifying languages into high, medium, and low-resource groups is important for our analysis.
Tables 2 and LABEL:ref:gen present important observations, which we outline below:
Extractive vs Abstractive Shows Same Pattern: The patterns shown by both the extractive and abstractive dataset results in Tables 2 and LABEL:ref:gen are consistent. This is due to the impact of the availability or lack of language-specific training data, and the quality of translation remaining the same for both types of QA tasks. Therefore, the insights and patterns observed in Table 2 also apply to Table LABEL:ref:gen.
Gemma2-Base performed best on our experiments: There is noticeable variation in the performance of different LLMs across languages. This can be attributed to the varying amounts of language-specific training data and the different sizes of the models. Despite this variation, the observation that translation-based methods work better for low-resource languages generally holds true across all LLMs.
Among the LLMs, Gemma2 is the clear winner across most languages and tasks. Llama-3 and Gemma perform similarly and slightly better than Bloom. The performance ranking of LLMs tends to follow their release order, suggesting that more training data improves their multilingual capabilities. Openhathi performs relatively poorly compared to other models, likely because it has only been trained on Hindi data and has limited exposure to other languages.
Additionally, Aya-101, Narvasa 2.0, and Mistral NeMo are superior instruction-tuned models. This can be attributed to their fine-tuning on Indic data, whereas other models perform well in their base variants but struggle during instruction fine-tuning. This difficulty is likely due to catastrophic forgetting, as the fine-tuning data for these models contains significantly less Indic data compared to their pre-training datasets. Aya-101 is fine-tuned over the base model mT5, which is an encoder-decoder model (non-LLM). Although it is 13B parameter model which is greater than other in terms of parameter count, then to its performance is way higher than other multilinguual instruction tunned models [5].
Languages Llam3.1 Instruct Gemma2-it Navarasa-2.0 Mistral-nemo Aya-101
Ext Gen Ext Gen Ext Gen Ext Gen Ext Gen
As 13.82 3.33 3.43 1.32 21.59 3.27 21.29 0.93 43.38 3.15
Bn 15.96 3.69 5.33 1.98 19.14 3.59 25.35 1.57 57.64 3.79
Gu 12.73 5.80 6.19 3.37 30.77 6.01 24.08 2.61 61.42 5.84
Hi 25.31 3.93 10.63 2.64 31.31 5.62 46.76 2.03 70.45 2.43
Kn 6.37 4.46 3.35 1.85 24.55 4.99 21.31 1.24 55.21 4.23
Ml 14.56 5.88 4.42 2.10 28.40 5.21 31.61 2.13 55.52 5.25
Mr 24.37 4.68 8.48 2.65 22.90 2.78 39.98 1.28 58.98 3.34
Od 14.56 5.04 2.99 0.95 20.75 3.12 0.26 0.60 48.82 3.23
Pa 10.25 4.21 3.59 1.73 25.84 5.88 32.04 2.42 64.19 6.01
Ta 13.95 6.23 5.75 2.97 18.67 6.57 28.04 2.40 58.27 7.07
Te 10.33 5.32 4.09 2.19 21.94 5.89 22.13 1.66 56.60 7.29
Table 5: Performance of various Large Language Models on Zero-Shot on Subset of INDIC QA BENCHMARK (MLQA, NQ open, Synthetic data). We report the average numbers (F1) across span-extraction datasets and (Rouge-L)abstractive question-answering datasets.
@@ A.2 Analysis
Table 6: Performance of various Large Language Models on Zero-Shot INDIC QA BENCHMARK. We report the average numbers (F1) across span-extraction datasets and (Rouge-L)abstractive question-answering datasets.
Languages Bloom Gemma Llama-3 Openhathi Gemma-2
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
Direct
Inference
Translate
As 13.86 16.19 21.70 16.08 19.23 24.14 1.14 5.38 33.01 35.60
Bn 17.84 16.07 21.15 16.69 19.14 23.96 1.49 5.46 38.33 35.11
Gu 13.27 18.59 24.90 25.12 20.27 29.55 4.07 7.15 40.10 44.86
Hi 21.69 19.29 34.37 19.72 41.53 30.02 11.83 8.28 44.46 43.18
Kn 15.63 17.86 24.47 16.31 20.39 25.84 0.31 6.24 35.31 37.98
Ml 19.22 17.96 25.20 17.31 22.80 28.07 0.68 6.74 38.35 41.54
Mr 15.12 18.89 23.96 17.46 36.12 27.54 2.07 6.74 41.89 41.10
Od 11.11 15.09 9.06 15.52 11.23 23.76 0.26 5.61 28.81 36.39
Pa 15.60 20.54 28.43 19.88 21.96 30.54 1.31 8.23 43.37 45.10
Ta 19.96 18.22 22.45 17.42 19.74 27.54 0.70 7.03 39.64 40.29
Te 16.07 17.79 23.93 17.72 13.59 25.37 0.53 6.38 34.97 39.24
En 27.41 27.13 34.69 8.55 52.17
Table 7: Performance of both Direct Inference and translate test inference of various Large Language Models on Zero-Shot Extractive INDIC QA BENCHMARK. We report the average numbers (F1) across span-extraction datasets.
@@
Figure 9: Base Model(Bloom) prediction on one of the instance of Abstractive QA dataset
Figure 10: Instruction tunned Model(Bloomz) prediction on one of the instance of Abstractive QA dataset.
@@
Figure 13: CHRF scores computed after translating the synthetic dataset from Hindi to asm and then back-translating from asm to Hindi.
Figure 14: CHRF and CHRF++ scores computed after translating the synthetic dataset from Hindi to ben and then back-translating from ben to Hindi.
Figure 15: CHRF and CHRF++ scores computed after translating the synthetic dataset from Hindi to guj and then back-translating from guj to Hindi.
Figure 16: CHRF and CHRF++ scores computed after translating the synthetic dataset from Hindi to ory and then back-translating from ory to Hindi.
Figure 17: CHRF and CHRF++ scores computed after translating the synthetic dataset from Hindi to kan and then back-translating from kan to Hindi.
Frequently Asked Questions
1) Why choose IndicTrans2 over other available translation models?
Answer: Gala et al. (2023) show that IndicTrans2 surpasses other models, such as NLLB and Google Translate, particularly for English to Hindi translation tasks. In our analysis, we also tested NLLB, but the CHRF scores for back-translated and source texts were lower than those achieved with the IndicTrans2 model.